HUBBLE SPACE TELESCOPE DAILY REPORT #2293 PERIOD COVERED: 0000Z (UTC) 01/22/99 - 0000Z (UTC) 01/25/99 Daily Status Report as of 025/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Five Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Seven Sets of WF/PC-2 7387 (T-Tauri Star Snapshot Survey) The WF/PC-2 was used to make observations of the stars DOAR-25, RNO-90, DOAR-24E, DOAR-21, SR-4, and GK-TAU as part of a T-Tauri class star survey. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Four Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.4 Completed Six Sets of STIS/CCD 8091 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude, non-scripted, parallel observations. During the third iteration of this proposal, the acquisition failure described in yesterday's report and HSTAR 6935 occurred, and the take data flag remained down. Otherwise, the observations were completed as planned, and no further anomalies were noted. 1.5 Completed Five Sets of STIS/CCD 8084 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted parallel high galactic latitude observations as part of POMS test proposal. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Two Sets of WF/PC-2 6967 (Blue Continuum Snapshots of 3CR Radio - Continued Galaxies) The WF/PC-2 was used to make observations of the radio galaxies 3C237.0, 3C1.0 and 3C465.0. Radio galaxies are an important class of extragalactic objects: they represent one of the most energetic astrophysical phenomena and they may be used as probes of their environments. They are unique probes of the early Universe. During the first iteration of this proposal, the acquisition failure described in yesterday's report and HSTAR 6935 occurred, and the take data flag remained down. Otherwise, the observations were completed as planned, and no further problems were reported. 1.7 Completed WF/PC-2 6346 (Snapshot Survey of the Ultraluminous IRAS Galaxy Sample-Part 1 High) The WF/PC-2 was used to make snapshot survey observations of the ultraluminous IRAS galaxies CL-ULIR15 and CL-ULIR235. These data will be searched for multiple cores, which imply a merger has occurred, and for starburst regions. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed WF/PC-2 6361 (High Resolution Imaging of X-Ray Selected AGNs) The WF/PC-2 was used to make high resolution observations of the known low red shifted x-ray active galactic nucleus (AGN) MS1030.2.2. These data will allow a study of the optical properties of an unbiased sample of AGN host galaxies. The observations were completed as scheduled, and no problems were reported. 1.9 Completed WF/PC-2 6354 (Measuring Luminosity Evolution in Z = 0.3 Field Galaxies) The WF/PC-2 was used to make snapshot observations of the galactic cluster Gal-Cluster-1 located at a distance of Z = 0.3. The targets were drawn from a much larger pool of galaxies for which the internal kinematics have been measured from an extensive ground based program. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed Two Sets of STIS/CCD/MA1/MA2 7761 (The Missing Mass of the Crab Nebula) The Space Telescope Imaging Spectrograph (CCD/MA1/MA2) was used to observe the Crab Nebula, along with its central pulsar. The observed mass in the nebula and pulsar is less than half of the expected initial stellar mass and the observed energy is only a few percent of the typical supernova energy. We propose to search for this fast shell by taking a spectrum of the Crab pulsar in the region of the C IV 1550 line. The observations were completed with no reported anomalies. 1.11 Completed Four Sets of WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed Two Sets of WF/PC-2 6651 (WLR Distances from Accurate Multi-Color WF/PC-2 Photometry of A & B Supergiants in M-31 and M-33) The WF/PC-2 was used to make multi-color photometric observations of NGC-224-40. This is to study A and B supergiant stars in M-31 and M-33 with a precision far exceeding that of existing ground-based estimates. The observations were completed as planned, and no problems were reported. 1.13 Completed Two Sets of STIS/CCD 8065 (Comparison of Protoplanetary Disks) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the protoplanetary disk around star HD31293. Multiple roll angles were used to minimize systematic errors and image artifacts. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed STIS/CCD 7361 (The Nature of Nuclear Activity in Nearby Galaxies) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the nuclear region of galaxy N2489. The high resolution spectra will be used to search for broad components in H-Alpha and for velocity gradients, indicative of gas rotation or other ordered motions. A wavelength calibration image was also taken. The loss of lock mentioned in 2.1 and HSTAR 6937 occurred about 45 minutes before the end of this approximate 4-hour, 15-minute proposal, affecting one observation. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.15 Completed STIS/CCD/MA2 7402 (Chromospheric Activity in Old Stars: Halo Subdwarfs' Variability) The Space Telescope Imaging Spectrograph (CCD and MA2) was used to make observations of the old sub-dwarf star HD94028 in order to study its chromospheric activity. All normal stars cool enough to have convection zones share a common minimum level of chromospheric activity which decreases with advancing spectral type; most show higher emission levels which vary with time. A wavelength calibration observation was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.16 Completed STIS/MA1/MA2 7651 (STIS MAMA Dispersion Solution Check) The Space Telescope Imaging Spectrograph (MA1 and MA2) was used to make MAMA dispersion solution checks. Wavelength dispersion solutions will be determined on a yearly basis as part of a long-term monitoring program. The observations were completed as scheduled, and no anomalies were reported. 1.17 Completed STIS/CCD/MA1/MA2 7400 (Evidence from Boron for Massive Star Evolution) The Space Telescope Imaging Telescope (CCD/MA1/MA2) was used to observe the B III and B II resonance lines at 2065.8 and 1362.5 Angstroms, respectively, in the main-sequence and slightly evolved blue star HD16592 using the HST STIS MAMA detectors and echelle gratings. The proposal completed with no reported problems. 1.18 Completed STIS/CCD/MA1/MA2 7653 (LSF Measure of the MAMA-Spectroscopic Modes) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make observations of HEN1357 to measure the LSF of MAMA spectroscopic modes with an external target. A wavelength calibration image was also taken. The loss of lock mentioned in 2.1 and HSTAR 6938 occurred about 30 minutes before the end of this approximate 2-hour proposal, affecting one observation. Otherwise, the observations were completed as planned, and no further anomalies were reported. 1.19 Completed STIS/CCD/WF/PC-2 7367 (Multi-Wavelength Imaging of Circumstellar Disks in the Orion Nebula) The Space Telescope Imaging Spectrograph (CCD) and the WF/PC-2 were used to make multi-wavelength imaging of circumstellar disks (NGC-1976) seen as silhouettes against the bright emission-line background in the Orion Nebula. With these data one should be able to examine the radial structure of the disks in great detail, and over a wide spatial extent. The observations were completed as planned, and no problems were reported. 1.20 Completed WF/PC-2 7622 (WF/PC-2 Cycle 7 Internal Monitor) The WF/PC-2 was used to make a set of Cycle 7 internal monitoring calibration measurements. A variety of internal exposures are obtained in order to provide a monitor of the integrity of the CCD camera electronics in both bays {gain 7 and gain 15}, a test for quantum efficiency in the CCDs, and a monitor for possible buildup of contaminants on the CCD windows. A bias calibration and an internal flat calibration image were also taken. The observations were executed as scheduled, and no anomalies were noted. 1.21 Completed Two Sets of WF/PC-2 8061 (POMS Test Proposal: WF-II Backup Parallel Archive Proposal) This is a POMS test proposal designed to simulate scientific plans. The proposal was executed as planned, and no anomalies were reported. 1.22 Completed WF/PC-2 6720 (The Evolution of Radio Loud Quasars II) The WF/PC-2 was used to make observations of the radio loud quasar MRC0040-20 as part of a mini-survey to study its evolution and morphology. The observations were executed as scheduled, and no anomalies were noted. 1.23 Completed WF/PC-2 6557 (WF/PC-2 Imaging of Intermediate Redshift Absorption Selected Galaxies) The WF/PC-2 was used to make observations of the galaxy Q0229-FIEL where a halo appears to be absorbing light from the background quasi-stellar object. The observations were completed as planned, and no problems were reported. 1.24 Completed STIS/CCD/MA1 7576 (Lyman-Alpha Absorption in Low-Z QSO's) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make observations of the Lyman-Alpha absorption in the low-Z quasar PKS0405-12. A luminous early-type galaxy 40" from the sight line is present at this redshift {projected distance = 78/h kpc}. A wavelength calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.25 Completed WF/PC-2 6357 (The Nuclear Morphology of Elliptical Galaxies) The WF/PC-2 was used to make images of the early type, large, elliptical galaxy ESO443-39 (taken from a list of 150 galaxies) in order to study its nuclear morphology. This survey will complement Illingworth's snapshot study of later Hubble types. The observations were completed as scheduled, and no anomalies were reported. 1.26 Completed STIS/CCD/MA1/MA2) 7395 (Chromospheric Activity in Old Stars: HD 106516) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make chromospheric observations of the old star HD106516. This is a repeat observation. This observation failed in Cycle 5 due to a failed guide star acquisition, and was not repeated at that time in accord with snapshot policy. A wave length calibration image was also taken. The observations were completed as scheduled, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 26 Successful : 26 Per HSTAR 6737, there was an approximate three minute loss of lock beginning at 023/115535Z. The proposal described in 1.14 was affected. Per HSTAR 6738, there was a loss of lock beginning at 023/205745Z. The proposal described in 1.18 was affected. Scheduled Reacquisitions : 22 Successful : 22 2.2 FHST Updates: Scheduled : 61 Successful : 61 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared six times. A TTR was written for a 2-minute, 13-second loss of recoverable 32k data due to equipment problems at STGT. DF-224 flight software version 10.1 was successfully installed at 023/1214Z. Version 10.2 was also successfully installed at 023/1250Z. ESTR-3 was reconditioned at 023/1553Z. ESTR-3 was then powered off. The STIS MCE-2 reset at 023/232641Z during a SAA crossing and while the low voltage was on. The STIS flight software error counter was reset at 024/000333Z per ROP NS-12. MAMA-2 was recovered via normal SMS commanding at 024/201854Z, the time of the next high voltage on commanding. As described in HSTAR 6942, there was a STIS EMC mechanism re-try at 024/114258Z. This produced a "routine" status buffer message. The STIS flight software error counter was reset at 024/115611Z per ROP NS-12. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH